OpenStudy (beccab003):

Can someone please help me with this question? I came up with a completely different answer. I will show you my steps down below. I will medal the person with the best answer. Thanks! Calculate the geometric mean of 5 and 125. 25 12.5 17.7 20

OpenStudy (beccab003):

5/125=125/x 5x=15625 x=3125

hartnn (hartnn):

if x is the geometric mean, then you must have \(\huge \dfrac{5}{x}=\dfrac{x}{125}\)

OpenStudy (johnweldon1993):

Oh, it should be \[\large \frac{5}{x} = \frac{x}{125}\]

OpenStudy (johnweldon1993):

Then cross multiply and solve for 'x'

hartnn (hartnn):

note that you will get 2 geometric means. only one is in the choices

OpenStudy (beccab003):

Okay so, x^2=625

OpenStudy (beccab003):

so x =25

hartnn (hartnn):

yes what are the 2 values of x ?

hartnn (hartnn):

x=25 and ?

hartnn (hartnn):

and 25 is the one which is in the choices :)

OpenStudy (beccab003):

How do I find a second value? That doesn't make sense. I thought there was only one value.

OpenStudy (johnweldon1993):

\[\large x^2 = 625\] \[\large x = \pm 25\] :)

hartnn (hartnn):

\(x^2 =a^2 \implies x =\pm a\)

hartnn (hartnn):

always remember that x^2 means its a quadratic equation, and x must have 2 values

hartnn (hartnn):

example : x^2 =100 gives x= 10 , -10 because (-10)^2 also = 100

OpenStudy (beccab003):

so it's -25?

hartnn (hartnn):

its 25, -25 and 25 is in the choices

OpenStudy (beccab003):

Okay thanks! You have been very helpful.

hartnn (hartnn):

welcome :)